Find the corresponding y values for x=-4, -3, -2 if y+x^+2x-15.
Answer by checkley77(12844)

y+x^[2]+2x-15 I'll assume you meant to have a [^2) power for the first x term?
When x=-4:
x(-4)=y+(-4^2)+2*-4-15
x(-4)=y+16-8-15
x(-4)=y-7 ans.
Where x=-3
x(-3)=y+(-3)^2+2*-3-15
x(-3)=y+9-6-15
x(-3)=y-12 ans.
Where x=-2
x(-2)=y(-2)^2+2*-2-15
x(-2)=y+4-4-15
x(-2)=y-15 ans.